# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v3

@@ IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the number $L$ and length $N$ of the chains, and contribute to reveal the inter-chain joint effect, to find optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12. It can be seen in Fig. 11(c1111, g1111, d1111) that in the absence of the cavity the number $L$ of chains has obvious effect on exciton transport of the three configurations SSH ST, TC-SSH ST, and TC-SSH TT for $L\leq 6$. For HE TT, although the input exciton current is large, the occupation number $p_{k,N}$ of the last column limits the outgoing exciton current, leading to that $L$ has not effect on exciton transport. In fact, because of the inefficient transport of excitons among chains for HE TT, the most of excitons locate at the upper edge sites [Fig. 11(b1)], which is similar to skin effect. On the contrary, because of the diffusion of excitons among multichains the inner occupation number $p_{\rm bulk-in}$ increase and the occupation number of the edge chains decrease with $L$ for HE ST as shown in Fig. 11(b1, e1, f1). Therefore, for HE ST the excitons are aggregated on the inside of the multichains with large $L$ [Fig. 11(e1)], thus it is perfect candidate applied to quantum battery in the six configurations. In the presence of the cavity the steady-state dynamics represent abundant phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinct periodic oscillation with $L$ for HE TT and HO-HE TT, in which the peaks (troughs) of $I_{\rm i,o}$ and $\eta$ correspond to the odd (even) numbers of $L$ for $N=10$ [Fig. 11(c2, g2, d2)]. The result is contrary to that of single chain. The difference between the steady-state dynamics of the odd and of the even number becomes small with $L$. In addition to, the exciton transport is strengthened by the cavity for HE ST and HE TT.
@@
There is an obvious character that the exciton transport efficiency $\eta$ behaves oscillation with the length $N$ of the chains in Fig. 12. We firstly discuss the situation without cavity in Fig. 12(a-d). The oscillation is weak for HO ST relative to the other configurations and depends on the parity of $L$ and $N$ for HO TT in the absence of cavity. Thus the diagonal inter-chain coupling induces difference between exciton transports of odd and of even $N$ in long HO TT multichains as shown in Fig. 12(b, d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for HE ST and HE TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing on the multichains for later. However, in the mixed long multichains of HO-HE because of the presence of HO chains the dimerization is suppressed, the oscillation is enhanced for HO-HE ST and HO-HE TT. For the two configurations the oscillations with $N$ are consistent for odd $L$ and are opposite for even $L$ each other, such as when $L=3$ and $N=16$ $\eta$ is trough for both HO-HE TT and HO-HE ST, while $L=4$ and $N=16$ $\eta$ is peak for HO-HE ST and is trough for HO-HE TT. When the cavity is considered the oscillation is significantly strengthened for the configurations containing HE chain, as shown in Fig. 12(e-h). For HE ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for HE ST and HE TT is consistent, which illustrates further that the oscillation is mainly caused by HE chain. Interestingly, the oscillation amplitude behaves inverse dependence on the parity of $L$ between HE TT and HO-HE ST, HO-HE TT, i.e. for the former it with odd $L$ is bigger than that with even $L$, and for the latter it with even $L$ is bigger than that with odd $L$. Furthermore, the dependence of the exciton transport efficiency for HO-HE ST and HO-HE TT on the parity of $N$ can be changed by $L$ and the cavity. For example, comparing $L=2$ and $L=3$ for HO-HE ST of $N=16$, $\eta$ corresponds respectively to peak and trough [Fig. 12(a, b)], as well as comparing Fig. 12(b) and 12(f) for HO-HE ST of $N=14$ and $L=2$ it corresponds respectively to trough and peak. We also note by comparing Fig. 12(b) and 12(f), Fig. 12(f) and 12(g) for HO-HE ST, that the period of the oscillation with $N$ also is varied by the cavity and $L$.
@@ V Summary
In summary, in order to reveal the mechanism of exciton transport in a 2D multichains system with a cavity, the spectra, the second order partial derivative of the Hopfield coefficients of cavity photons with respect to the inter-chain coupling strength and the dimerization parameter, the von Neumann entropy and the steady-state dynamics of exciton are obtained by numerically diagonalizing the Hamiltonian and solving the Lindblad quantum master equation, in which six inter-TLS coupling configurations are chosen as HO ST, HO TT, HE ST, HE TT, HO-HE ST and HO-HE TT based on the HO and HE chain, as well as the square and triangle type of inter-chain coupling. The effects of the inter-chain coupling strength, the dimerization parameter, the detuning of the driving frequency to the TLS, the number of chain, the length of chain, and the coupling configurations on them are investigated. The results show that in the absence of cavity the exciton transport efficiency is decided by the distribution of population of exciton on the whole multichains. Besides, when the cavity is considered the occupation number of the cavity photons depend significantly on the cavity-dressed energy level crossings and anticrossings of the nearly-zero-energy modes, at which the coherent excitation and LZ transition between the multichains and the cavity occur. Therefore, the exciton transport efficiency and currents via the polariton channel are changed at the locations of the crossings and anticrossings. We can conclude that the exciton transport dynamics are controlled by the crossings and anticrossings. Besides, the locations of the crossings and anticrossings is accurately determined by calculating von Neumann entropy. We also discover that the second order partial derivative of the Hopfield coefficients of cavity photons is a good physical quantity to describe the locations of the crossings and anticrossings, and the change of photonic occupation number. The spectra show that the locations of the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter. Particularly, the exciton transport efficiency and currents are symmetric about the inter-chain coupling strength $\zeta=0$ in the multi-chain system without cavity, but the symmetry can be broken by the cavity. In the space of the dimerization parameter the presence of the cavity enhances exciton transport for HE ST, HO-HE ST and HO-HE TT with even chains, but suppresses it for HO-HE ST and HO-HE TT with odd chains when $\delta>0.36$. Interestingly, when the cavity drive is simultaneously exerted, the exciton transport is not limited by the occupation number of site $11$ and is strengthened for HE ST and HO-HE ST with odd chains in the range of the negative dimerization parameter. Further, in the space of the TLS transition frequency detuning and the inter-chain coupling strength we discover that the exciton transport is suppressed for HO ST and HO TT and is strengthened for the other configurations by the cavity at the range corresponding to the crossings and anticrossings, in which the polaritons act as vital role and the effects are enhanced by the cavity drive. In the space of the TLS transition frequency detuning and the dimerization parameter, the steady-state dynamics depend on the coupling configurations, the driven methods, and the cavity, especially in the zone of the fully dimerized limit. Finally, it can be found that the exciton transport is obviously related to the parity of the number and length of the chains with cavity, in which the exciton transport efficiency behaves oscillation with them. The oscillation is related to the configurations. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like the topological phase transition of the system.
@@ VI Acknowledgements
This work is supported by the Academic Ability Promotion Foundation for Young Scholars of Northwest Normal University in China under the Grant No. NWNU-LKQN2024-10, National Natural Science Foundation of China Regional Fund under Grant No. 12164042, National Natural Science Foundation of China under Grant No. 12104374, National Natural Science Foundation of China under Grant No. 12065022.

### v4

@@ IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the chain number $L$ and length $N$, which reveal the inter-chain joint effect, and give optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12.
@@
Fig. 11(c1111, g1111, d1111) shows that in the absence of the cavity the exciton transport is mightily dependent on chain number $L$ for the three configurations HE ST, HO-HE ST, and HO-HE TT when $L\leq 6$, and is opposite for the others. For HE TT, although the input exciton current $I_{i}$ is large, the little $p_{N}$ limits the outgoing exciton current $I_{o}$, leading to that exciton transport is independent of $L$. In fact, the inefficient exciton transport among chains lead to excitons localizing at the upper edge sites [Fig. 11(b1)], which is similar to skin effect. Contrary to the skin effect, for HE ST because of the diffusion of excitons among chains the inner occupation number $p_{\rm bulk-in}$ increase and that of the edge decrease with $L$ as shown in Fig. 11(b1, e1, f1). For large $L$ the excitons collect at the inside of the multichains [Fig. 11(e1)], thus HE ST is a potential candidate applied to quantum battery in the six configurations. In the presence of the cavity the steady-state dynamics present characteristic phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinctly odd-even oscillation with $L$ for HE TT and HO-HE TT, in which the peaks (troughs) of $I_{\rm i,o}$ and $\eta$ correspond to the odd (even) $L$ when $N=10$ [Fig. 11(c2, g2, d2)]. The difference between the steady-state dynamics of the odd and even $L$ becomes small with $L$. In addition, it can be seen that only for HE ST and HE TT the exciton transport is strengthened by the cavity when $N=10$.
@@
There is an obvious character in Fig. 12 that the exciton transport efficiency $\eta$ also behaves odd-even oscillation with the chain length $N$. When the cavity is absent, the oscillation is weak for HO ST relative to the other configurations and depends simultaneously on the odevity of $L$ and $N$ for HO TT [Fig. 12(a-d)]. Thus the diagonal inter-chain coupling induces the difference between the exciton transports of odd and even $N$ in long HO TT multichains as shown in Fig. 12(b, d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for HE ST and HE TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing on the multichains for the later. However, in the mixed long multichains of HO-HE, because HO chains suppress the dimerization, the oscillation is enhanced for HO-HE ST and HO-HE TT. For the two configurations the oscillations with $N$ are consistent for odd $L$ and are opposite for even $L$ each other, such as when $L=3$ and $N=16$ $\eta$ is trough for both HO-HE TT and HO-HE ST, while $L=4$ and $N=16$ $\eta$ is peak for HO-HE ST and is trough for HO-HE TT. When the cavity is considered the oscillation is significantly strengthened for the configurations containing HE chain, as shown in Fig. 12(e-h). For HE ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for HE ST and HE TT is consistent, which illustrates further that the oscillation is mainly caused by HE chain. Interestingly, the oscillation amplitude behaves inverse dependence on the odevity of $L$ between HE TT and HO-HE ST, HO-HE TT. Concretely, for the former the amplitude with odd $L$ is bigger than that with even $L$, and for the latter it is contrary. Furthermore, it can be discovered that for HO-HE ST and HO-HE TT the odd-even oscillation of $\eta$ with $N$ can be adjusted by $L$ and the cavity. For example, the oscillations of $L=3$ and $L=4$ for HO-HE ST are out of phase [Fig. 12(a, b)]. $\eta$ without [Fig. 12(b)] and with [Fig. 12(f)] cavity for HO-HE ST of $N=14$ and $L=2$ are respectively trough and peak. Comparing Fig. 12(b) and 12(f), Fig. 12(f) and 12(g) for HO-HE ST, the period of the oscillation also is varied by the cavity and $L$.
@@ V Summary
In summary, in order to reveal the exciton transport mechanism of a 2D multichains system within a cavity, we investigate systematically the spectra through numerically diagonalizing the Hamiltonian, the second order partial derivative of photon Hopfield coefficients with respect to the inter-chain coupling strength and the dimerization parameter, the von Neumann entropy and the steady-state dynamics of exciton via solving the Lindblad quantum master equation, in which six inter-TLS coupling configurations are considered as HO ST, HO TT, HE ST, HE TT, HO-HE ST and HO-HE TT based on the HO and HE chain, as well as the square and triangular inter-chain coupling type. The results show that the exciton transport can be adjusted by the inter-chain coupling strength, the dimerization parameter, the detuning of the driving frequency, the chain number and length, the coupling configurations, and the cavity. In the absence of cavity the exciton transport is decided by the distribution of exciton on whole multichains, in which the efficient transport need the excitons spread all over the multichains. Further, when the cavity is considered it is governed by the competition between the polariton and multichain channel. When the excitations are mainly distributed on whole multichains, the multichains dominate the transport. Otherwise, the cavity occupies the most excitations, the polaritons act as the main channel, meanwhile the limited exciton occupations of the multichain terminals are necessary. The exchange condition of the excitation between the multichains and cavity is the key which channel is primary. The number of polaritons depend on the photon occupation number, which is relate to the cavity-dressed nearly-zero-energy crossings and anticrossings. At the crossings and anticrossings the coherent excitation and LZ transition between the multichains and cavity occur, respectively. Therefore, the exciton transport is controlled by the crossings and anticrossings, whose locations can be accurately determined by von Neumann entropy. Besides, we discover that the second order partial derivative of the photon Hopfield coefficient serves as robust indicator for identifying the crossings and anticrossings, and the change of photon occupation number. The spectra show that the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter. In the space of inter-chain coupling strength, ST outperforms TT for the exciton transport, but the optimal ranges differ for HO and HE chains. In the space of the dimerization parameter, even-length ST is more beneficial to exciton transport than TT under TLS drive, but odd-length ST is under TLS+cavity drive. Interestingly, when the cavity drive is exerted, the exciton transport is not limited by the occupation number of site (1,1) and is strengthened for HE ST and HO-HE ST with odd chains in the range of the negative dimerization parameter. Further, we discover that the steady-state dynamics can be adjusted by the TLS transition frequency detuning. Finally, it can be found that the exciton transport currents and efficiency present distinctly odd-even oscillation with chain length and number, which depends on the coupling configuration and the cavity. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like the topological phase transition of the system.
@@ VI Acknowledgements
This work is supported by the Academic Ability Promotion Foundation for Young Scholars of Northwest Normal University in China under the Grant No. NWNU-LKQN2024-10, National Natural Science Foundation of China under Grant No. 12164042, National Natural Science Foundation of China under Grant No. 12104374, National Natural Science Foundation of China under Grant No. 12065022.